If tan mƟ = tan nƟ, then the different  values of Ɵ will be

Grade:12th pass

1 Answers

Krish Kumar
3 years ago
If m = n, then, tan(mθ) = tan(nθ) has infinite solutions.
 
But
 
If m ≠ n, then,
mθ = kπ + nθ where k is any integer.
(because the solutions of tan(x) = tanα are given by x = kπ + α where k is any integer.
So, θ(m-n) = kπ
Hence, θ = kπ/(m-n)